OpenStudy (dimples97):

What percent of 60 is 12? Need help on solving this please, I need to show my work

sam (.sam.):

\[\frac{x \%}{100 \%} \times 60 = 12\]

sam (.sam.):

Solve for x

OpenStudy (dimples97):

divide 60 ?

sam (.sam.):

It works, divide both sides by 60 \[\frac{x \%}{100 \%} \times 60 = 12 \\ \\ \frac{x \%}{100 \%} = \frac{12}{60} \\ \\ \frac{x \%}{100 \%} = \frac{1}{5} \]

OpenStudy (dimples97):

got 0.2 its same as 1/2?

sam (.sam.):

Cancel the percent \[ \frac{x \cancel \%}{100 \cancel \%} = \frac{1}{5} \\ \\ \frac{x}{100}=\frac{1}{5}\] Multiply both

sam (.sam.):

Multiply both sides by 100 and you're done

OpenStudy (dimples97):

Ohhh i see, thankyou!!!

sam (.sam.):

Yw :)

OpenStudy (wolf1728):

To be specific, 12 is 20% of 60

OpenStudy (wolf1728):

12/60 = x*100% 1200 / 60 = 20%